# SplineSketch: Even More Accurate Quantiles with Error Guarantees

### v1

@@ 3. Description of SplineSketch
We provide a description of our algorithm, omitting implementation details, explained in Sec. 4. In fact, there are many possible implementations of the sketch that still satisfy the worst-case guarantee. Our sketch works for any numerical input, i.e., the input items are integers or floating-point numbers of bounded precision. The theoretical analysis of the worst-case error is deferred to Appendix  B.
For simplicity, we first present the sketch assuming there are no items of high frequency, namely appearing at least $n/k$ times in the input after inserting any number $n$ of items into SplineSketch of size $k$. We show how to remove this assumption in Sec. 3.3.
@@
Fix an integer parameter $k\geq 6$ (as required in the analysis). The sketch primarily consists of $k$ distinct thresholds $\tau_{1}<\tau_{2}<\cdots<\tau_{k}$ and $k$ counters $b_{1},\ldots,b_{k}$, where $b_{i}$ is our estimate of the number of items that lie in the interval $(\tau_{i-1},\tau_{i}]$; we define $\tau_{0}=-\infty$. We call these intervals buckets and refer to a bucket by the index of its right threshold, so $(\tau_{i-1},\tau_{i}]$ is called ”bucket $i$” or the $i$-th bucket. We also frequently refer to $b_{i}$ as the size of bucket $i$ and we define the length of bucket $i$ as $\ell_{i}:=\tau_{i}-\tau_{i-1}$. Throughout the execution of the algorithm $\tau_{1}$ and $\tau_{k}$ are always set to be the minimum and the maximum item in the stream, respectively. This makes the first bucket special as its counter only stores the frequency of $\tau_{1}$.
We also maintain a buffer of size $O(k)$ that we use for faster processing of incoming items, i.e., low average update time, and one auxiliary bit-array of length $k$ for “protected thresholds” (Sec. 3.1). After processing the whole input, the buffer is merged into the buckets and the auxiliary bit-array can be discarded. We maintain the two invariants: All buckets counters plus the number of items in the buffer sum up to $n$ and that no bucket is empty.
@@
When a new element arrives, we first put it in the buffer. When the buffer gets full, we call the consolidate method (Sec. 3.1) that may change bucket thresholds and merges the buffer into buckets. The only exception is the first time when the buffer gets full. In that case, we simply sort it, pick every $n/k$-th item, thus selecting $k$ items, and set the counters $b_{i}$ to exactly the number of items in the corresponding buckets. Since no item has frequency at least $n/k$, these selected items are distinct.
@@
For a query at a threshold $\tau_{i}$, we return the sum of the counters $\sum_{j=1}^{i}b_{i}$. For $x$ strictly inside the $i$-th bucket, i.e. $\tau_{i-1}<x<\tau_{i}$, we use the PCHIP interpolation to estimate the rank of $x$; note that this works no matter whether $x$ appeared on input or not. We remark that computing this interpolation requires only the knowledge of $\tau_{i-2},\cdots,\tau_{i+1}$ together with corresponding prefix sums of bucket counters. The interpolation does not extrapolate, i.e., for $x<\tau_{1}$, it returns 0, and for $x>\tau_{k}$, it returns $\sum_{j=1}^{k}b_{i}$. Since one has to find the right bucket by binary search, one query takes time $O(\log k)$ provided that the prefix sums are precomputed.
@@ 3.1. Consolidating buckets
The main procedure of the sketch is the consolidate method. The method first counts the new number of items in each bucket if we were to merge the buffer into the buckets. Then, it keeps changing the buckets based on the projected new values of the counters, possibly removing some of the thresholds and creating new ones. After those changes, the buffer gets merged into the new buckets, meaning we again count for each bucket the number of items currently in the buffer that fall into that bucket and update the counters accordingly. Finally, the buffer gets emptied. The crux of the consolidate method is how to modify the thresholds, which we describe in the remainder of this subsection.
@@
We have two basic operations that change bucket thresholds, join and split. A join takes two neighboring buckets $(\tau_{i-1},\tau_{i}]$ and $(\tau_{i},\tau_{i+1}]$ and turns them into a single bucket $(\tau_{i-1},\tau_{i+1}]$ with its counter set to $b_{i}+b_{i+1}$. Note that after joining, buckets are (implicitly) renumbered. That is, we actually set new thresholds $\tau^{\prime}$ and counters $b^{\prime}$ as $\tau^{\prime}_{i}=\tau_{i+1}$, $b^{\prime}_{i}=b_{i}+b_{i+1}$, and for any $j>i$, $\tau^{\prime}_{j}=\tau_{j+1}$ and $b^{\prime}_{j}=b_{j+1}$. The buckets up to $\tau_{i-1}$ do not change.
When splitting bucket $(\tau_{i-1},\tau_{i}]$, we replace it by two buckets $(\tau_{i-1},\frac{\tau_{i-1}+\tau_{i}}{2}]$ and $(\frac{\tau_{i-1}+\tau_{i}}{2},\tau_{i}]$. The bucket counters are set as follows: We use the query method to get estimated ranks $\hat{r}$ at $\tau_{i-1},\frac{\tau_{i-1}+\tau_{i}}{2},\tau_{i}$. We then set the counters for the two new buckets to $\hat{r}(\frac{\tau_{i-1}+\tau_{i}}{2})-\hat{r}(\tau_{i-1})$ and $\hat{r}(\tau_{i})-\hat{r}(\frac{\tau_{i-1}+\tau_{i}}{2})$, respectively. Other buckets are again implicitly renumbered.
Observe that every split increases and every join decreases the number of buckets by one. Thus, we accompany every split with joining two buckets, which ensures there are always $k$ buckets (unless a change of $k$ is desired). We argue in Appendix B.2 that this can always be done in a way that preserves worst-case guarantees.
@@
If the buffer contains a new minimum $\tau^{\prime}_{1}$ (an item smaller than the previous minimum $\tau_{1}$), we add a new bucket $(-\infty,\tau^{\prime}_{1}]$ before all of the existing buckets, set its counter to the frequency of $\tau^{\prime}_{1}$ in the buffer, and we also increase the counter of the bucket $(\tau^{\prime}_{1},\tau_{1}]$, where $\tau_{1}$ is the previous minimum, by the number of buffer items in this interval. A new maximum $\tau^{\prime}_{k}$ is handled similarly, by adding a new bucket $(\tau_{k},\tau^{\prime}_{k}]$ and setting its counter to the number of buffer items in this interval. For each of these at most two new buckets, we perform a join of two buckets. The rules for selecting which pairs buckets to join are described in the following paragraphs. Note that these at most two new buckets may be split during the consolidation that created them.
Maintaining bounds on bucket counters and heuristic error.
We maintain two invariants, the first of which is used to prove worst-case bounds while the second ensures good performance in practice.
In order to obtain low worst-case uniform error we need to ensure that no bucket has too large counter. To this end, we require that all bucket counters are bounded by $O(n/k)$, specifically
@@
for some constant $C_{b}>1$ that will be determined in the analysis in Section B.3.
Secondly, we define a heuristic error for each bucket, which is intended to capture “hard” parts of the input, where the input distribution changes a lot, namely, where the derivative of the empirical PDF is large. For such buckets, the PCHIP interpolation may have larger error than for buckets with low derivative of the PDF. Since the empirical PDF in the bucket $i$ is $b_{i}/\ell_{i}$, we use the following heuristic error for bucket $i$, which approximates the second derivative of the empirical CDF, normalized by length squared:
@@
For the edge cases $i=1$ and $i=k$, we make the following assumptions: $b_{0}=b_{k+1}=0$, $\ell_{0}=\ell_{1}$, and $\ell_{k+1}=\ell_{k}$; that is, we add “virtual” buckets 00 and $k+1$ with zero counters and a length equal to the length of the adjacent (real) bucket.
We now describe how we use the bucket bound (1) and heuristic error (2) when performing the splits and joins. Each time we split a bucket $i$, we choose a pair of adjacent buckets to join in order to keep the total number of buckets the same. This pair is selected so that after joining it would yield a bucket with the lowest heuristic error, of all joinable pairs, defined as follows, which are not overlapping bucket $i$.
@@
We say that a pair of adjacent buckets is joinable if both of the following conditions hold:
@@
    the bucket resulting from the join would not violate the bucket size bound (1) or be relatively close to it, specifically it would have counter at most $0.75\cdot C_{b}\cdot n/k$, and
@@
    any of the two buckets were not subject to a split recently; specifically, threshold $\tau_{i}$ is not protected as explained below.
We perform splits as follows. First, we take all buckets that exceed the size bound (1). We split them one by one, always performing a join satisfying conditions of Definition 1 for each split; in Lemma 2 we show that at least one pair of adjacent buckets can be joined. We also join at most two further pairs of joinable buckets if the buffer contains a new minimum or maximum, each creating a new bucket. Note that a bucket may need to be split repeatedly during one call of the consolidate method, namely, if it exceeds the bucket bound (1) substantially.
After all these necessary splits and joins, we consider a bucket $i$ with a high heuristic error (2), and we split it if its error is more than $\gamma$ times greater than the lowest heuristic error of a bucket resulting from joining a joinable pair, for a parameter $\gamma>1$ (say, $\gamma=1.5$). Furthermore, we do not perform a split of a bucket $i$ due to the heuristic error if there are less than $k/3+2$ pairs of adjacent buckets that can be joined, which is needed to prove that there is a joinable pair of buckets; see Lemma 2. Finally, if the counter of bucket $i$ is substantially below the bound, say, $b_{i}\leq(C_{b}/100)\cdot n/k$, we also do not split bucket $i$, thus avoiding to split almost empty buckets. We summarize the conditions for splitting:
@@
We repeat finding a splittable bucket $i$ and a joinable pair of buckets not overlapping bucket $i$, performing the split and join as long as there are splittable buckets. Lemma 2 shows that if there is a splittable bucket $i$, then there is a joinable pair of buckets not overlapping bucket $i$.
Protected thresholds and epochs.
The above description yields a sketch with high accuracy in practice, but contrived instance may cause a large error, much larger than $n/k$. Basically, the issue is that one bucket may be split and joined repeatedly, while accumulating error; in the extreme case, one bucket may undergo $\Omega(k)$ splits and joins, and each time, the error at the threshold increases by $\Omega(n/k)$, leading to $\Omega(n)$ error. With the small overhead of $k$ bits when processing the input, we show how to avoid such adversarial repeated splits and joins at a single location by protecting thresholds of buckets that were recently created by a split.
Specifically, we divide the input into epochs. The first epoch has length $\Theta(k)$ and each successive epoch starting at time $t$ has length $0.25\cdot t$. When we split a bucket $(\tau_{i-1},\tau_{i}]$, we mark all three thresholds $\tau_{i-1},\tau^{\prime}_{i},\tau_{i}$ as protected, where $\tau^{\prime}_{i}=(\tau_{i-1}+\tau_{i})/2$ is the new threshold. We store a bit vector of length $k$, with the $i$-th bit set to 1 when the $i$-th threshold is protected. When a new epoch starts, we reset all thresholds to the default unprotected state by zeroing the bit vector. As specified in Definition 1, we never join buckets that have a protected threshold between them.
Note that we may still split a single bucket multiple times during one epoch, or even during one call of the consolidate method (namely, if the bucket receives a large number of items). Similarly, a single bucket may be joined multiple times during one consolidation, but a combination of splitting and joining is not allowed – bucket which was created by a split cannot be joined with another bucket in the same epoch.
Resizing sketch during consolidate.

### v2

@@ 3. Description of SplineSketch
We provide a description of our algorithm, omitting implementation details, explained in Sec. 4. In fact, there are many possible implementations of the sketch that still satisfy the worst-case guarantee. Our sketch works for any numerical input, i.e., items are integers or floating-point numbers. We first present the sketch assuming there are no items of high frequency, namely appearing more than (n1)/(k1)(n-1)/(k-1) times in the input after inserting any number $n$ of items into SplineSketch of size $k$. We show how to remove this assumption in Sec. 3.3.
@@
Fix an integer parameter $k\geq 6$ (as required in the analysis). The sketch primarily consists of $k$ distinct thresholds $\tau_{1}<\tau_{2}<\cdots<\tau_{k}$ and $k$ counters $b_{1},\ldots,b_{k}$, where $b_{i}$ is our estimate of the number of items that lie in the interval $(\tau_{i-1},\tau_{i}]$; we define $\tau_{0}=$-$\infty$. We call these intervals buckets and refer to a bucket by the index of its right threshold, so $(\tau_{i-1},\tau_{i}]$ is called ”bucket ii” or the $i$-th bucket. We also frequently refer to $b_{i}$ as the size of bucket $i$ and we define the length of bucket $i$ as $\ell_{i}:=\tau_{i}-\tau_{i-1}$.
Throughout the execution of the algorithm $\tau_{1}$ and $\tau_{k}$ are always set to be the minimum and the maximum item in the stream, respectively. This makes the first bucket special as its counter only stores the frequency of $\tau_{1}$.
We also maintain a buffer of size $O(k)$ that we use for faster processing of incoming items, i.e., low average update time, and one auxiliary bit-array of length $k$ for “protected thresholds” (Sec. 3.1). After processing the whole input, the buffer is merged into the buckets and the auxiliary bit-array can be discarded. We maintain two invariants: All bucket counters plus the number of items in the buffer sum up to $n$ and no bucket is empty.
@@
When a new element arrives, we first put it in the buffer. When the buffer gets full, we execute the consolidate method (Sec. 3.1) that may change bucket thresholds and merges the buffer into buckets. The only exception is the first time when the buffer gets full. In that case, we simply sort it, pick the first and then every $\approx(n-1)/(k-1)$-th item, selecting $k$ items as the initial thresholds, and set counters $b_{i}$ to exactly the number of items in the corresponding buckets. Since no item has frequency greater than (n1)/(k1)(n-1)/(k-1), these selected items are distinct and every bucket is nonempty. For the pseudocode of the bucket initialization, see Algorithm 1.
@@
For a rank query at a threshold $\tau_{i}$, we return the sum of the counters up to the $i$-th, i.e., $\sum_{j=1}^{i}b_{i}$. For $x$ strictly inside the $i$-th bucket, i.e. $\tau_{i-1}<x<\tau_{i}$, we use an interpolation method to estimate the rank of $x$; note that this works no matter whether $x$ appeared on input or not. The purpose of using a suitable interpolation is to get much better error in practice, while not affecting the worst-case bounds (see Sec. 3.3).
Specifically, we use a monotone variant of the Piecewise Cubic Hermite Interpolating Polynomial (PCHIP) (Fritsch and Butland, 1984; Fritsch and Carlson, 1980), which is an interpolation method that constructs a continuous function by fitting cubic polynomials between data points. Unlike standard cubic spline interpolation, PCHIP adjusts the first derivatives at each data point to preserve the shape and monotonicity of the original data. For an introduction in greater detail, including the formulas for the interpolation, see (Wong, 2020). Since the ranks of the input items are non-decreasing, preservation of monotonicity makes PCHIP suitable for our application.
We remark that computing this interpolation requires only the knowledge of $\tau_{i-2},\cdots,\tau_{i+1}$ together with corresponding prefix sums of bucket counters. Therefore, provided that we maintain the prefix sums, the computation of the interpolation takes $O(1)$ time. The interpolation does not extrapolate, i.e., for $x<\tau_{1}$, it returns 0, and for $x>\tau_{k}$, it returns $\sum_{j=1}^{k}b_{i}$. Since one has to find the right bucket by binary search, one query takes time $O(\log k)$ provided that the prefix sums are precomputed.
In principle, our approach can be used with any interpolation method (ideally one that preserves monotonicity). One popular alternative would be to use kernel smoothers. However, this particular method would introduce additional error on the thresholds, which would be harder to control. We leave the exploration of alternative interpolation methods for future work. For an extensive treatment of different interpolation methods, see (Fan and Gijbels, 1996).
@@ 3.1. Consolidating buckets
The main procedure of the sketch is Consolidate. The method first merges the buffer into the current buckets, by counting the new number of items in each bucket, and updates the counters accordingly. Then, it keeps changing the buckets based on the new values of the counters, possibly removing some of the thresholds and creating new ones. Finally, the buffer gets emptied. The crux of the consolidate method is how to modify the thresholds, which we describe in the remainder of this subsection. For the pseudocode of Consolidate, see Algorithm 2.
Figure 3. Illustration of joining a pair of buckets and splitting one bucket (depicted similarly as in Fig. 2). The join removes threshold $\tau_{j}$ and merges (sums up) the two counters of the joined buckets. The split takes the $i$-th bucket $(\tau_{i},\tau_{i+1}]$ and inserts a new threshold $\tau^{\prime}_{i+1}=(\tau_{i}+\tau_{i+1})/2$ in the middle, setting the counters of the two resulting buckets $(\tau_{i},\tau^{\prime}_{i+1}]$ and $(\tau^{\prime}_{i+1},\tau_{i+1}]$ using the interpolation (solid black curve). The thresholds $\tau_{i},\tau^{\prime}_{i+1}$, and $\tau_{i+1}$ get protected from joining, as depicted by the locks, and this protection is removed at the end of the epoch.
@@
We have two basic operations that change bucket thresholds, join and split (Fig. 3). A join takes two neighboring buckets $(\tau_{i-1},\tau_{i}]$ and $(\tau_{i},\tau_{i+1}]$ and turns them into a single bucket $(\tau_{i-1},\tau_{i+1}]$ with its counter set to $b_{i}+b_{i+1}$. Note that after joining, buckets are (implicitly) renumbered. That is, we actually set new thresholds $\tau^{\prime}$ and counters $b^{\prime}$ as $\tau^{\prime}_{i}=\tau_{i+1}$, $b^{\prime}_{i}=b_{i}+b_{i+1}$, and for any $j>i$, $\tau^{\prime}_{j}=\tau_{j+1}$ and $b^{\prime}_{j}=b_{j+1}$. The buckets up to $\tau_{i-1}$ do not change.
When splitting bucket $(\tau_{i-1},\tau_{i}]$ for $i>1$, we replace it by two buckets $(\tau_{i-1},\frac{\tau_{i-1}+\tau_{i}}{2}]$ and $(\frac{\tau_{i-1}+\tau_{i}}{2},\tau_{i}]$. The bucket counters are set as follows: We use the query method to get estimated ranks $\hat{r}$ at $\tau_{i-1},\frac{\tau_{i-1}+\tau_{i}}{2},\tau_{i}$. We then set the counters for the two new buckets to $\hat{r}(\frac{\tau_{i-1}+\tau_{i}}{2})-\hat{r}(\tau_{i-1})$ and $\hat{r}(\tau_{i})-\hat{r}(\frac{\tau_{i-1}+\tau_{i}}{2})$, respectively. Other buckets are again implicitly renumbered.
Observe that every split increases and every join decreases the number of buckets by one. Thus, after every split we join some other pair of buckets, which ensures there are always $k$ buckets (unless a change of $k$ is desired). The crux of the algorithm is in selecting which buckets to split and join, which we describe below.
Protected thresholds and epochs.
After splitting a bucket, the estimated rank, and therefore the estimation error, at its right threshold stays the same. However, the error at the newly created threshold may be bigger, in the worst case by the size of the original bucket counter. If we repeatedly join and split the newly created buckets with their neighbors, we might let the error increase too much. With the small overhead of $k$ bits when processing the input, we show how to avoid too frequent splits and joins at a single location and prevent the associated error accumulation. To this end we introduce the notion of protected thresholds.
Specifically, we divide the input into epochs. The first epoch has length $\Theta(k)$ and each successive epoch starting at time $t$ has length $0.25\cdot t$. When we split a bucket $(\tau_{i-1},\tau_{i}]$, we mark all three thresholds $\tau_{i-1},\tau^{\prime}_{i},\tau_{i}$ as protected, where $\tau^{\prime}_{i}=(\tau_{i-1}+\tau_{i})/2$ is the new threshold. We store a bit vector of length $k$, with the $i$-th bit set to 1 when the $i$-th threshold is protected. When a new epoch starts, we reset all thresholds to the default unprotected state by zeroing the bit vector. As specified later in Def. 1, we never join buckets that have a protected threshold between them.
Note that we may still split a single bucket multiple times during one epoch, or even during one call of Consolidate, namely, if the bucket receives many new items. Similarly, a single bucket may be joined multiple times during one consolidation, but a combination of splitting and then joining is not allowed – a bucket which was created by a split cannot be joined with another bucket in the same epoch.
@@
If the buffer contains a new minimum $\tau^{\prime}_{1}$ (an item smaller than the previous minimum $\tau_{1}$), we add a new bucket $(-\infty,\tau^{\prime}_{1}]$ before all of the existing buckets, set its counter to the frequency of $\tau^{\prime}_{1}$ in the buffer. We also increase the counter of the bucket $(\tau^{\prime}_{1},\tau_{1}]$ by the number of buffer items in this interval. A new maximum $\tau^{\prime}_{k}$ is handled similarly, by adding a new bucket $(\tau_{k},\tau^{\prime}_{k}]$ and setting its counter to the number of buffer items in this interval. For each of these at most two new buckets, we perform a join of another pair, selected according to the rules described below. Note that these at most two new buckets may be later split during Consolidate in which they were created.
Maintaining bounds on bucket counters and small heuristic error.
We maintain two invariants, the first of which is used to prove worst-case bounds while the second ensures high accuracy in practice.
In order to obtain a low worst-case uniform error, we need to ensure that no bucket has a too large counter. To this end, we require that all bucket counters are bounded by $O(n/k)$, specifically
@@
for some constant $C_{b}>1$ that will be determined in the analysis in App. B.3.
Secondly, we define a heuristic error for each bucket, which is intended to capture “hard” parts of the input, where the input distribution changes a lot, namely, where the second derivative of the empirical CDF is large. For such buckets, the PCHIP interpolation may have larger error than for buckets with a low second derivative. Since the empirical PDF (i.e., first derivative of the CDF) in the bucket $i$ is $b_{i}/\ell_{i}$, we use the following heuristic error for bucket $i>1$, which approximates the second derivative of the empirical CDF, normalized by length squared:
@@
For the edge cases $i=2$ and $i=k$, we make the following assumptions: $b_{k+1}=0$, $\ell_{1}=\ell_{2}$, and $\ell_{k+1}=\ell_{k}$; that is, we add a “virtual” buckets $k+1$ with zero counter and a length equal to the length of the adjacent (real) bucket, and we set the length of bucket 1 to the length of bucket 2.
We now describe how we use the bucket bound (1) and heuristic error (2) when performing the splits and joins. Each time we split a bucket $i$, we choose a pair of adjacent buckets that do not contain buckets created from splitting $i$ and we join that chosen pair in order to keep the total number of buckets the same. The pair is selected so that the bucket resulting from joining has the lowest heuristic error of all joinable pairs, defined as follows. In a nutshell, the bucket resulting from the join should not violate the bucket size bound (1) or be relatively close to it. Moreover, none of the two buckets should be a subject to a split recently, which we implement using the protected thresholds.
@@
We say that a pair $i,i+1$ of adjacent buckets is joinable if both of the following conditions hold:
@@
    the counter of the bucket resulting from the join satisfies $b_{i}+b_{i+1}\leq 0.75\cdot C_{b}\cdot n/k$, and
@@
    threshold $\tau_{i}$ is not protected.
We perform splits as follows. First, we take all buckets that exceed the size bound (1). We split them one by one, always performing a join satisfying conditions of Definition 1 for each split; in Lemma B.2 we show that at least one pair of adjacent buckets can be joined. We also join at most two further pairs of joinable buckets if the buffer contains a new minimum or maximum, each creating a new bucket. Note that a bucket may need to be split repeatedly during one call of Consolidate, namely, if it exceeds the bucket bound (1) substantially.
After all these necessary splits and joins, we consider a bucket $i$ with a high heuristic error (2), and we split it if its error is more than $\gamma$ times greater than the lowest heuristic error of a bucket resulting from joining a joinable pair, for a parameter $\gamma>1$ (say, $\gamma=1.5$). Furthermore, we do not perform a split of a bucket $i$ due to the heuristic error if there are less than $k/3+2$ pairs of adjacent buckets that can be joined, which is needed to prove that there is a joinable pair of buckets; see Lemma B.2. Finally, if the counter of bucket $i$ is substantially below the bound, say, $b_{i}\leq(C_{b}/100)\cdot n/k$, we also do not split bucket $i$, avoiding splitting almost empty buckets. We summarize the conditions for splitting:
@@
We repeat finding a splittable bucket $i$ and a joinable pair of buckets not overlapping bucket $i$, performing the split and join as long as there are splittable buckets. Since every split increases the number of protected thresholds, we know that this process terminates after at most $O(k)$ iterations. In Appendix B.2, we show it is also well-defined.
Lemma 3 (Lemma 2).
SplineSketch always contains a joinable pair of adjacent buckets, according to Definition 1.
Input : Buffer of size $n$, parameter $k$ (number of buckets)
Output : Initialized $k$ buckets with thresholds $\tau_{1},\dots,\tau_{k}$ and counters $b_{1},\dots,b_{k}$
1 Sort all $n$ items in the buffer in ascending order;
2 For $i=0,\dots,k-1$, pick the item at index $\min\{\lceil i\cdot(n-1)/(k-1)\rfloor,n-1\}$ from the sorted buffer to form $k$ thresholds $\tau_{1},\dots,\tau_{k}$;
3 For each bucket $(\tau_{i-1},\tau_{i}]$, set the counter $b_{i}$ to the exact number of items in the buffer that fall into this interval;
4 Return the initialized buckets with thresholds $\tau_{1},\dots,\tau_{k}$ and counters $b_{1},\dots,b_{k}$;
Algorithm 1 Buckets Initialization
Input : $k$ buckets and buffer of new items
Output : Updated buckets and protection bit vector
1
2if $n\geq t$ then
3   Reset protection bit vector, and set $t:=1.25\cdot t$;
4
5Incorporate buffer items into each bucket’s counter, and empty the buffer;
6 if new minimum or maximum found in the buffer then
7   Add a new boundary bucket for minimum/maximum;
8   For each new bucket, find a pair of joinable buckets $j,j+1$, and join this pair;
9
10while exists a splittable bucket $(\tau_{i-1},\tau_{i}]$, first processing those exceeding the bucket bound (1) do
11   Split bucket $i$ into $(\tau_{i-1},\tau^{\prime}]$ and $(\tau^{\prime},\tau_{i}]$ for $\tau^{\prime}=(\tau_{i}+\tau_{i-1})/2$;
12   Mark thresholds $\tau_{i-1},\tau^{\prime},\tau_{i}$ as protected;
13   Join a pair of joinable buckets $j,j+1$ (Def. 1) that does not overlap with $(\tau_{i-1},\tau^{\prime}]$ and $(\tau^{\prime},\tau_{i}]$ and has the lowest heuristic error;
14
Algorithm 2 Consolidate
Resizing the sketch during Consolidate.